A person walks 20.0° north of east for 3.06 km. How far would she have to walk due north and due east to arrive at the same location?

____ km (due north)
____ km (due east)
Answer #1

sin (20) = O/H

O = sin (20) * 3.06
O = 1.04658 km north

cos (20) = A/H

A= cos(20) * 3.06
A = 2.8754 km east

total = 3.922 km

=============================================================================